OpenStudy (freemap):

The function t(h) = 2h – 10 gives the outdoor temperature, in degrees Fahrenheit, h hours after midnight. What is the meaning of the y-intercept? A. The temperature was –10°F at midnight. B. The temperature was 5°F at midnight. C. The temperature was 0°F at 5 a.m. D. The temperature was 0°F at 10 a.m.

OpenStudy (bloomlocke367):

Well, can you point out the y-intercept in the equation?

OpenStudy (freemap):

-10?

OpenStudy (texaschic101):

you mean the y int is -10

OpenStudy (bloomlocke367):

yes my bad XD my dad walked over and I got sidetracked while typing XD

OpenStudy (bloomlocke367):

When an equation is in the form of y=mx+b (which is what you have) m=slope and b=y-intercept, so yes, the y-intercept is -10 c:

OpenStudy (texaschic101):

look at it this way...the h = hours after midnight. so if you sub in 0 for h...0 hrs after midnight....you would get the temp at midnight. t(h) = 2h - 10 y = 2h - 10 y = 2(0) - 10 y = -10 so that is the temp at midnight

OpenStudy (freemap):

Makes sense leaving you with A for the correct answer

OpenStudy (texaschic101):

yep :)

OpenStudy (bloomlocke367):

Yep! :) @texaschic101 explained it better than what I could XD

OpenStudy (texaschic101):

don't know about that...I probably just confused the crap out of him...lol

OpenStudy (bloomlocke367):

haha XD

OpenStudy (freemap):

@texaschic101 has always been my tutor. She's one of the best. Great at helping confused people. Solves the most complicated problems. Thanks @texaschic101 I love to see when you come online

OpenStudy (texaschic101):

awww......* blushes *....thx

OpenStudy (freemap):

You earned it!

OpenStudy (texaschic101):

I try.....I am giving you an imaginary hug right now

OpenStudy (freemap):

awww so sweet, such a wonderful person, just because you know your studies, I like how you're not like rude.

OpenStudy (texaschic101):

rude won't get you anywhere.

OpenStudy (freemap):

I'm not going to take your day, so you have a wonderful rest of your day, hopefully I won't be on here again, I know you have other things.

OpenStudy (texaschic101):

and just think...before long, you will be helping others with problems such as these :) Have a good day and stay out of trouble...lol

OpenStudy (freemap):

lol I don't have a problem with trouble, Its easy for me to stay in line